### v5

# Matthew Blackwell   Jacob R. Brown   Sophie Hill Kosuke Imai   Teppei Yamamoto

Abstract
Conditioning on variables affected by treatment can induce post-treatment bias when estimating causal effects. Although this suggests that researchers should measure potential moderators before administering the treatment in an experiment, doing so may also bias causal effect estimation if the covariate measurement primes respondents to react differently to the treatment. This paper formally analyzes this trade-off between post-treatment and priming biases in three experimental designs that vary when moderators are measured: pre-treatment, post-treatment, or a randomized choice between the two. We derive nonparametric bounds for interactions between the treatment and the moderator under each design and show how to use substantive assumptions to narrow these bounds. These bounds allow researchers to assess the sensitivity of their empirical findings to priming and post-treatment bias. We then apply the proposed methodology to a survey experiment on electoral messaging.
Keywords: bounds, interactions, heterogeneous effects, measurement, moderation, sensitivity analysis
1 Introduction
Ascertaining heterogeneous treatment effects is an integral part of many survey experiments. Researchers are often interested in how treatment effects vary across respondents with different characteristics. For example, we may be interested both in how implicit versus explicit racial cues affect support for a particular policy but also in how those effects differ by levels of racial resentment (Valentino, Hutchings and White, 2002). Or we may want to know whether the effect of land-based electoral appeals might depend on the voters’ sense of land security (Horowitz and Klaus, 2020). These questions of effect heterogeneity allow researchers to explore potential causal mechanisms and design more targeted and effective future treatments.
To examine such treatment effect heterogeneity, we must measure the relevant covariates, such as racial resentment or land security in the aforementioned examples, at some point during the survey experiment. The question of when we measure these moderators, however, is a source of methodological debate. On the one hand, a long tradition in political science has recognized the potential priming bias of a pre-test design, where covariates are measured prior to treatment (e.g., Transue, 2007; Morris, Carranza and Fox, 2008; Klar, 2013; Klar, Leeper and Robison, Forthcoming; Schiff, Montagnes and Peskowitz, 2022). For example, asking a respondent about their party identification might lead them to evaluate the treatment in a more partisan or political light, resulting in biased causal effect estimates. Several studies have documented priming effects from a range of different covariates (see Klar, Leeper and Robison, Forthcoming, for a review). Some find that certain priming effects can last for weeks (Chong and Druckman, 2010).
On the other hand, the practice of measuring moderators after treatment, what we call a post-test design, has come under scrutiny due to the possibility for post-treatment bias (Rosenbaum, 1984; Acharya, Blackwell and Sen, 2016; Montgomery, Nyhan and Torres, 2018). In particular, if covariates are affected by the treatment, then conditioning on those covariates—as required when assessing effect heterogeneity—can bias the estimation of conditional average treatment effect and any interactions that compare such effects. In our empirical application, the key moderating variable of land insecurity is a subjective, perceived measure and thus potentially affected by the framing of a political appeal around land rights. Though treatment is unlikely to affect measurements of many moderators like basic demographics, researchers investigating manipulable moderators face a dilemma about when to measure these covariates when designing an experiment.
In this paper, we apply the nonparametric identification and bounding approach of Manski (1995) and Balke and Pearl (1997) to formally analyze the trade-off between priming and post-treatment biases under different experimental designs, and propose principled ways to analyze data (see Imai and Yamamoto, 2010, for a similar analysis of measurement error).
We begin by deriving nonparametric bounds to show that neither the pre-test nor post-test design provides much information about conditional average treatment effects or interactions without additional assumptions. Next, we show how three potentially plausible assumptions can narrow the bounds. The first is priming monotoncity, which assumes that whether the moderator is measure pre- or post-treatment only affects the outcome in one direction. The second is moderator monotonicity, which assumes that measuring the covariates after treatment can move the value of that moderator only in one direction. The third assumption is stable moderator under control, whereby the covariate under the control condition cannot be affected by the timing of treatment. None of these assumptions can point identify the interaction between the treatment and a moderator, but they can substantially narrow the bounds and sometimes be informative about the sign of such an interaction.
To further sharpen our inference, we generalize the two standard experimental designs and consider a randomized placement design, where the experimenter randomly assigns respondents to either the pre-test or post-test design. We demonstrate how to estimate nonparametric bounds in this context and how to incorporate assumptions that connect the pre-test and post-test arms to further narrow the bounds. In our supplemental materials, we also develop a parametric Bayesian approach to incorporate pre-treatment covariates in the analysis to sharpen our inferences and quantify estimation uncertainty.
We also derive sensitivity analysis procedures for all three designs. In these analyses, we vary the proportion of respondents whose outcomes or moderators are affected by when the moderator is measured and assess how the bounds change as a function of this sensitivity parameter. This procedure allows researchers and readers to gauge the credibility of an estimated interaction in light of a more nuanced set of assumptions, rather than the blunt instruments of the monotonicity and stability.
Several recent studies have empirically explored the trade-off between priming and post-treatment bias. Albertson and Jessee (2022) find that the effect of moderator placement has little effect on the estimated interaction in a question-wording experiment. Furthermore, they find no evidence for an average effect of treatment on the moderator, potentially reducing concerns about post-treatment bias. Sheagley and Clifford (2023) compared several experiments when the moderators were measured just prior to treatment or in a prior survey wave. The authors found that estimated effects and interactions were similar across these conditions and concluded that priming bias may not be a widespread concern for experimental studies in political science.
Both of these papers present compelling evidence for the specific experiments conducted in their empirical assessments, but as Sheagley and Clifford (2023) warn us, “we should be cautious in generalizing [these] findings to the wide variety of studies run by political scientists.” Our approach, on the other hand, provides a general methodological toolkit that can be applied to any experimental design and allows researchers to include substantive assumptions to tailor the framework to their applications.
The rest of the paper proceeds as follows. We first introduce a motivating empirical example of how land insecurity moderates the effectiveness of land-based appeals by politicians from Horowitz and Klaus (2020). We next describe the notation and basic assumptions of the pre-test and post-test designs. We then derive the sharp nonparametric bounds and sensitivity analyses for the pre-test, post-test, and randomized placement designs. Next, we apply the proposed methods to the empirical example. Finally, we conclude by suggesting directions for future research.
2 Motivating Example
We illustrate the trade-off between the pre- and post-treatment measurement of a moderator using a survey experiment conducted by Horowitz and Klaus (2020). This study investigates the effectiveness of land-based appeals to increase a politician’s electoral support in Kenya’s Rift Valley. We focus on two randomly assigned conditions: a control condition in which participants heard a generic campaign speech with no direct reference to the land issue, and a treatment condition that additionally referenced the land issue. [^1] To preserve statistical power, we focus on a treatment group that combines two conditions: one stating the land issues is their top priority and one that additionally reference an ethnic grievance. The outcome is the participant’s reported likelihood of supporting the candidate, which we dichotomize (likely to support the candidate vs. not) to illustrate our proposed methods.
One of the main hypotheses tested by Horowitz and Klaus (2020) is whether individuals who are personally experiencing land insecurity are more responsive to land-based appeals by politicians. While the evidence from the full sample is inconclusive, among respondents belonging to the “insider” ethnic group (Kalenjins), there is a positive and statistically significant interaction between the treatment condition and a dummy variable for land insecurity, in line with the authors’ expectations.
Both types of bias could be of concern in this context, as land insecurity is measured by asking respondents to rate the security of their land rights. Asking this question before treatment may raise the salience of the land issue and thus attenuate any treatment effect of hearing a land-based appeal by a politician, a form of priming bias. Conversely, if the experimenter asks respondents about land security after treatment, then their responses may be affected by the content of the speech, yielding post-treatment bias.
To address these concerns, Horowitz and Klaus (2020) use (what we call) the randomized placement design, in which questions relating to the respondent’s land rights were randomly assigned to occur before or after the treatment. Figure 1 shows the “naïve” estimates of the treatment-moderator interaction, which are all positive and substantively large, implying that the effect of a land-based appeal on electoral support is 25–50 percentage points larger for land-insecure respondents compared land-secure respondents. However, these estimates only reach conventional levels of statistical significance with the increased statistical power of the combined pre/post sample.
Figure 1: Estimates of treatment-moderator interaction using pre-test, post-test, and combined data from Horowitz and Klaus (2020), with and without covariates (age, gender, education, and closeness to own ethnic group). Error bars indicate 95% confidence intervals.
What can we learn about the true interaction from these estimates? We show in the following sections that the observed data often tells us very little about interactions without strong assumption. In Section 5, we return to this empirical example to illustrate how researchers can assess the possible impact of priming bias and post-treatment bias on their substantive findings using our proposed methods.
3 Experimental Designs and Causal Quantities of Interest
We now lay out the formal notation of our setting and describe the causal quantities of interest. Let $T_{i}$ represent the binary treatment variable for unit $i\in\{1,\ldots,n\}$, indicating a certain experimental manipulation received by the unit. We use $Y_{i}$ to denote a binary outcome of interest, and $M_{i}$ to represent an observed binary moderator of interest. [^2] We focus on the binary setting to ease exposition, though many of the methods could be extended to more general cases. Our goal is to develop a methodology that can be used to understand how the average treatment effect of $T_{i}$ on $Y_{i}$ varies as a function of $M_{i}$.
In this paper, we consider three experimental designs for this goal: the pre-test design, the post-test design, and the randomized placement design. These three designs differ in the timing of measurement for the potential moderators. In the pre-test design, the experimenter measures the moderator, $M_{i}$, before treatment assignment, ensuring that these measurements are unaffected by treatment. In the post-test design, the experimenter measures moderators after treatment assignment. Let $Z_{i}$ be an indicator for measuring $M_{i}$ before ($Z_{i}=0$) or after treatment ($Z_{i}=1$).
We will analyze these experimental designs using the potential outcomes framework (e.g., Holland, 1986). Let $Y_{i}(t,z)$ represent the potential outcome with respect to the treatment status $t$ and the timing of covariate measurement, $z$. In our empirical application, for example, $Y_{i}(1,1)$ is whether respondent $i$ would support the hypothetical candidate if they were given the speech with referencing the land issue (that is, treatment, $t=1$) and we measured the land security moderator after treatment ($z=1$). We make the consistency assumption for the potential outcomes, such that $Y_{i}=Y_{i}(T_{i},Z_{i})$.
Our goal is to estimate treatment effects on the outcome free of priming bias, which requires us to measure moderators after treatment. To this end, we define the “true” potential outcomes of interest as $Y^{*}_{i}(t)\equiv Y_{i}(t,1)$ and let $Y_{i}(t)\equiv Y_{i}(t,0)$ to be a possibly mismeasured proxy observed under the pre-test design. Priming bias can occur when the pre-test and post-test measurements of the outcome differ, $Y_{i}(t)\neq Y^{*}_{i}(t)$.
Similarly, the measures of the moderator, self-reported land security, can be affected by the when it is measured and the land grievance framing treatment. Let $M_{i}(t,z)$ be the potential value of the moderator that would be observed for unit $i$ when the treatment is set to $t$ and the variable is measured in design $z$. Under the pre-test design, the treatment cannot affect the moderator, so we refer to $M_{i}(0,0)=M_{i}(1,0)\equiv M^{*}_{i}$ as the “true” moderator for unit $i$. On the other hand, under the post-test design, we let $M_{i}(t)\equiv M_{i}(t,1)$ to be potentially mismeasured proxy for that true moderator. Under those designs, the treatment can affect respondents’ moderator, so that $M_{i}(0)\neq M_{i}(1)$, which can lead to post-treatment bias.
We can explore heterogeneous treatment effects by estimating the following quantities of interest,
$\tau(m)$$\equiv$ $\mathbb{E}(Y_{i}^{*}(1)-Y^{*}_{i}(0)\mid M^{*}_{i}=m),$ (1)
$\delta$$\equiv$ $\tau(1)-\tau(0),$ (2)
where $m\in\{0,1\}$. The first quantity characterizes the post-test conditional average treatment effect (CATE) as a function of the pre-test value of the moderator. This CATE is the effect of land-based appeals on support for a politician (when unprimed) conditional on a level of true land insecurity. The second quantity of interest compares the CATE between two subpopulations with different levels of the pre-test moderator, which is often called an interaction. These two quantities of interest formalize the dilemma about the choice of pre-test versus post-test experimental designs. Identifying $\tau(m)$ requires observing the true potential outcomes of interest ($Y^{*}_{i}(t)$) and true moderator ($M^{*}_{i}$) for the same unit, but this can never occur because the former requires $Z_{i}=1$ and the latter requires $Z_{i}=0$.
While the pre-test design allows us to observe the true moderator, it may suffer from priming bias because asking questions about the moderator might change the causal effect of the treatment by cueing respondents. Thus, under the pre-test design, neither $\tau(m)$ nor $\delta$ nor even the ATE, $\mathbb{E}(Y^{*}_{i}(1)-Y^{*}_{i}(0))$, is identified. In contrast, under the post-test design, the ATE can be estimated without bias, and yet this design may result in post-test bias for $\tau(m)$ and $\delta$ when the treatment affects the moderator.
4 Nonparametric Analysis of the Three Experimental Designs
We now show what we can learn from each of the three possible experimental designs. Generally speaking, none of the three designs are informative about the sign of the interaction effect without further assumptions. For each design, we derive sharp bounds for the interaction and show how to narrow these bounds with additional substantive assumptions. We also develop a sensitivity analysis procedure that allows researchers to vary the strengths of such assumptions and assess their implications.
4.1 Pre-test Design
We first investigate the identifying power of the pre-test design, where we observe $(Y_{i},M_{i},T_{i})$ among the units for whom $Z_{i}=0$. We first formalize the randomization of treatment in this setting.
Assumption 1 (Pre-test Randomization).
$\{Y_{i}(t,z),M_{i}(t,1)\}\ \mbox{$\perp\!\!\!\perp$}\ T_{i}$$\mid$ $M^{*}_{i}=m,Z_{i}=0,$ (3)
for $t=0,1$, $m\in\{0,1\}$.
This assumption allows for the possibility of conditioning our randomization on the pre-treatment moderator, though this assumption also holds when randomization is unconditional.
By computing the difference in the average observed outcomes between different treatment groups, researchers can identify the following quantity,
$\tau_{pre}(m)$$\equiv$ $\mathbb{E}(Y_{i}\mid T_{i}=1,M_{i}=m,Z_{i}=0)-\mathbb{E}(Y_{i}\mid T_{i}=0,M_{i}=m,Z_{i}=0)$ (4)
$=$ $\mathbb{E}(Y_{i}(1)-Y_{i}(0)\mid M^{*}_{i}=m),$
where the equality follows from Assumption 1 and the consistency assumption. Equation (4) does not generally equal the true CATE, $\tau(m)$, since the conditional distribution of $Y_{i}(t)$ given $M^{*}_{i}=m$ may differ from that of $Y^{*}_{i}(t)$ given $M^{*}_{i}=m$, since asking questions about the moderator before measuring outcome may distort subsequent responses. Thus, $\tau_{pre}(m)$ is the CATE of the land grievance treatment on the primed support for the candidate. Similarly, $\delta$ is not generally equal to $\delta_{pre}\equiv\tau_{pre}(1)-\tau_{pre}(0)$. The key problem here is that the pre-test design gives us information about the correct values of the moderator $M^{*}_{i}$, but provides no information about the outcomes we want to investigate, $Y^{*}_{i}(t)$.
With no restrictions on the priming bias, we cannot rule out extreme possibilities, such as all respondents changing their values of $Y_{i}$ in response to the priming. This could occur if asking about land security caused all supporters to oppose the hypothetical candidate and vice versa. While perhaps unlikely, this scenario is possible under the randomization assumption alone. Thus, the nonparametric bounds under the pre-test design remain identical to the logical bounds,
$\tau(m)$$\in$ [1, 1],11[-1,\ 1],[ - 1 , 1 ] , (5)
$\delta$$\in$ [2, 2].22[-2,\ 2].[ - 2 , 2 ] . (6)
In other words, the pre-test design is completely uninformative about the causal effects of interest unless one is willing to make additional assumptions about the joint distribution of $Y^{*}_{i}(t)$ and $Y_{i}(t)$.
We can derive an expression for the amount of priming bias as
$\tau_{pre}(m)-\tau(m)$ =$=\left\{\mathbb{P}(Y^{*}_{i}(1)=0,Y_{i}(1)=1\mid M^{*}_{i}=m)-\mathbb{P}(Y^{*}_{i}(0)=0,Y_{i}(0)=1\mid M^{*}_{i}=m)\right\}$ (7)
$\quad-\left\{\mathbb{P}(Y^{*}_{i}(1)=1,Y_{i}(1)=0\mid M^{*}_{i}=m)-\mathbb{P}(Y^{*}_{i}(0)=1,Y_{i}(0)=0\mid M^{*}_{i}=m)\right\},$
where we have suppressed the conditioning on $Z_{i}=0$ to reduce notational burden. The first term of this bias for the CATE is the effect of treatment on probability of being positively primed (moving from $Y^{*}_{i}=0$ to $Y_{i}=1$) and the second term is the effect of treatment on being negatively primed (moving from $Y^{*}_{i}=1$ to $Y_{i}=0$). Positively primed individuals in our empirical application would be those that would become supportive of the candidate only when asked about land security prior to treatment. We can see that this bias is unidentifiable because it depends on the joint distribution of the unprimed $Y_{i}^{*}(t)$ and primed $Y_{i}(t)$ outcomes, but we never observe both of these variables for the same unit.
Some scholars, such as Sheagley and Clifford (2023), have shown empirically that $\delta$ and $\delta_{pre}$ appear close to each other in several experiments where the moderator was measured in a prior survey wave (and so was less prone to priming bias). Should we conclude from this that future studies will not suffer from priming bias? Perhaps, but we should properly view this as an additional strong assumption rather than a result implied by the design of the experiment. In particular, we would have to assume that either there is no priming at all, treatment does not affect priming, or that the treatment effects on positively and negatively primed individuals are exactly equal to each other and cancel out. Below, we will see how to incorporate assumptions about limited priming that would allow a sensitivity analysis for pretest designs.
4.1.1 Narrowing the Pre-test Bounds under Additional Assumptions
What assumptions can we place on the pre-test design to narrow the bounds? Any such assumption would have to place restrictions on the joint distribution of the primed and unprimed outcomes. A common set of assumptions for this type of setting would be a monotone treatment response assumption for the priming effect (Manski, 1997). In particular, we consider a priming monotonicity assumption, which states that the effect of asking the moderator before treatment can only move the outcome in a single direction.
Assumption 2 (Priming Monotonicity).
$Y_{i}(t)\geq Y^{*}_{i}(t)$ for all $t=0,1$.
The assumption implies that the effect of moving from post-test ($Z_{i}=1$) to pre-test ($Z_{i}=0$), which we call the priming effect, can only increase the outcome. In our application, this assumption implies that asking about land rights before reading the campaign speech treatment increases support for the hypothetical candidate. As stated, this assumption holds across levels of treatment, though it is possible to assume the reverse direction ($Y_{i}(t)\leq Y^{*}_{i}(t)$) or even to have a different effect direction for each level of treatment.
Proposition 1 (Pre-test Sharp Bounds).
Let $P_{tzm}=\mathbb{P}(Y_{i}=1\mid T_{i}=t,Z_{i}=z,M_{i}=m)$. Under Assumptions 1 and 2, we have the following sharp bounds:
$\tau(m)\in\left\{-P_{00m},\;P_{10m}\right\},$ (8)
and
$\delta\in\left\{-P_{100}-P_{001},\;P_{101}+P_{000}\right\}.$ (9)
Priming monotoncity narrows the bounds compared to the uninformative randomization bounds, but the bounds still will always contain zero. Thus, without further assumptions, the pre-test design cannot rule out no CATE for any level of the moderator nor can it rule out no interaction.
4.1.2 Sensitivity Analysis for the Pre-test Design
Given the empirical results such as Sheagley and Clifford (2023) pointing to a limited role of priming in some experimental settings, we now develop a sensitivity analysis for pre-test designs to determine how sensitive results are to deviations from the “no priming bias” assumption. In particular, we constrain the proportion of respondents primed, or more precisely, the proportion of respondents whose value of $Y^{*}_{i}(t)$ is different from $Y_{i}(t)$. Formally, we state this restriction as
$\Pr(Y^{*}_{i}(t)=1,Y_{i}(t)=0\mid M^{*}_{i}=m)+\Pr(Y^{*}_{i}(t)=0,Y_{i}(t)=1\mid M^{*}_{i}=m)\ \leq\ \theta$ (10)
for all $t,m\in\{0,1\}$. Note that if $\theta=0$, then we have $Y^{*}_{i}(t)=Y_{i}(t)$ for all $i$, and the data under the pre-test design alone identify the quantities of interest. By increasing the value of $\theta$, a sensitivity analysis gradually restricts the severity of the priming effects in the empirical setting.
Proposition 2 (Pre-test Sharp Bounds under Restricted Priming Effects).
Suppose that Assumptions 1 and 2, and restriction (10) hold. Then, we have the following sharp bounds:
$\tau(m)\in\left\{\tau_{pre}(m)-\theta,\tau_{pre}(m)+\theta\right\},\qquad\delta\in\left\{\delta_{pre}-2\theta,\delta_{pre}+2\theta\right\}.$
This proposition says that if the proportion of primed respondents in each treatment arm is no larger than $\theta$, then we can bound the true interaction with a $4\theta$ wide interval around the naive pre-test interaction (for the CATE, the interval width is exactly half, i.e., $2\theta$). These bounds can be informative (that is, they exclude 0) if $|\delta_{pre}|>2\theta$.
4.2 Post-test Design
Next, we consider the post-test design, where we observe $(Y_{i},M_{i},T_{i})$ among the units for whom $Z_{i}=1$. The randomization of the treatment implies the following ignorability assumption.
Assumption 3 (Post-Test Randomization).
$\{Y_{i}(t,z),M_{i}(t),M^{*}_{i}\}\ \mbox{$\perp\!\!\!\perp$}\ T_{i}\mid Z_{i}=1,$
for $t=0,1$.
How informative is Assumption 3 alone about the CATEs and interaction without making any other assumption? The standard CATE estimator would be unbiased for
$\tau_{post}(m)$$\equiv$ $P_{11m}-P_{01m}=\mathbb{E}[Y^{*}_{i}(1)\mid M_{i}(1)=m]-\mathbb{E}[Y^{*}_{i}(0)\mid M_{i}(0)=m],$ (11)
where the equality follows from Assumption 3 and the fact that $\Pr(Z_{i}=1)=1$ for all $i$.
Under the post-test design, we observe the true potential outcome of interest $Y_{i}^{\ast}(t)$, and yet the moderator may be observed with error, i.e., $M_{i}(t)\neq M_{i}^{\ast}$. Similar to the case of the pre-test design, therefore, $\tau_{post}(m)$ does not generally equal $\tau(m)$ because the conditional distribution of $Y^{*}_{i}(t)$ given $M_{i}(t)=m$ may differ from that of $Y^{*}_{i}(t)$ given $M^{*}_{i}=m$. In fact, $\tau_{post}(m)$ may not equal $\tau(m)$ even if the effects of treatment and moderator measurement timing do not change the marginal distribution of the moderator, i.e., $\Pr(M_{i}(0)=m)=\Pr(M_{i}(1)=m)=\mathbb{P}(M_{i}^{*}=m)$ for all $m\in\mathcal{M}$. Restricting the marginal distribution of the moderator does not help because effect heterogeneity is a function of the joint distribution of the counterfactual outcomes and moderators. Thus, under the post-test design, neither $\tau(m)$ nor $\delta$ is nonparametrically identified.
Despite the lack of point identification, the design can contain some information about the causal quantities of interest. The following proposition derives the sharp (i.e., shortest possible) randomization-only bounds under the post-test design.
Proposition 3 (No-assumption Post-test Sharp Bounds).
Let $P_{t}=\mathbb{P}(Y_{i}=1\mid T_{i}=t,Z_{i}=1)$. Under Assumptions 3, we have $\delta\in[\delta_{L},\delta_{U}]$, where
$\delta_{L}$ $=-1-\max\left\{\min\left(\frac{1-P_{0}}{P_{1}},\frac{1-P_{1}}{P_{0}}\right),\min\left(\frac{P_{1}}{1-P_{0}},\frac{P_{0}}{1-P_{1}}\right)\right\},$ (12)
$\delta_{U}$ $=1+\min\left\{\max\left(\frac{1-P_{0}}{P_{1}},\frac{1-P_{1}}{P_{0}}\right),\max\left(\frac{P_{1}}{1-P_{0}},\frac{P_{0}}{1-P_{1}}\right)\right\}.$
Furthermore, these bounds are sharp.
The proof of this result is given in Section A.2 and rely on a standard linear programming approaches often used in bounding causal quantities (e.g., Balke and Pearl, 1997). Unfortunately, these bounds are often quite wide in practice. In particular, the bounds can never be narrower than $[-1,1]$ since the post-test data are completely uninformative about the true moderator. As a result, the sharp bounds under the post-test design can be quite wide and sometimes cover the entire possible range, $[-2,2]$, for $\delta$. In sum, without additional assumptions, the post-test design can only provide limited information about the causal quantities of interest.
4.2.1 Narrowing the Post-test Bounds under Additional Assumptions
While the bounds only using the randomization can be too wide to be useful in practice, we may be willing to entertain other assumptions on the causal structure that will narrow the bounds. We rely on a principal stratification approach in which we stratify the units according to how their moderator values react to treatment and moderator measurement timing (Frangakis and Rubin, 2002). Let $\mu_{s}(t)=\mathbb{P}[Y^{*}_{i}(t)=1\mid S_{i}=s]$, where $S_{i}$ represents the principal strata defined by the moderator, $\{M_{i}(1),M_{i}(0),M^{*}_{i}\}$. Without making any assumption, $S_{i}$ can take any of the $2^{3}$ values in
$\mathcal{S}=\{111,011,101,001,110,010,100,000\}.$
Let $\rho_{s}=\Pr[S_{i}=s]$ be the probability of a unit falling into one of the strata, such that $\sum_{s\in\mathcal{S}}\rho_{s}=1$. Finally, we denote the marginal probability of the true pre-test moderator as $Q_{*}=\mathbb{P}(M^{*}_{i}=1)$.
The first assumption we consider is that the effect of post-treatment measurement of the moderator has a monotonic effect on the moderator for every unit.
Assumption 4 (Moderator Monotonicity).
$M_{i}(t)\geq M^{*}_{i}$ for all $t=0,1$.
In the context of the motivating example, this assumption requires no unit to have (say) lower land security values if we ask about it after the respondent reads the politician’s speech rather than before. While weaker than assuming there is no measurement error in the post-test moderator, the plausibility of this assumption will depend on the experimental context. In the post-test design, we cannot verify this assumption because we never observe $M_{i}^{*}$. The assumption rules out several possible principal strata, ensuring that $S_{i}$ can only take one of the following values: $111,110,010,100,$ or $000$. [^3] While we present a positive version of monotonicity for both treatment levels, it is possible to derive bounds under a negative version of the assumption or with differing directions for each treatment condition. We now present the sharp bounds under this assumption.
Proposition 4 (Post-test Sharp Bounds under Monotonicity).
Let $Q_{tz}=\Pr(M_{i}=1\mid T_{i}=t,Z_{i}=z)$. Under Assumptions 3 and 4, we have sharp bounds $\delta\in[\delta_{L2},\delta_{U2}]$, where
$\delta_{L2}$ $=\frac{P_{111}Q_{11}-P_{011}Q_{01}}{Q_{*}}-\frac{P_{110}(1-Q_{11})-P_{010}(1-Q_{01})}{1-Q_{*}}$
$\qquad+\frac{\max\{P_{011}Q_{01}-Q_{*},0\}-\max\{P_{111}Q_{11},Q_{11}-Q_{*}\}}{Q_{*}(1-Q_{*})},$
$\delta_{U2}$ $=\frac{P_{111}Q_{11}-P_{011}Q_{01}}{Q_{*}}-\frac{P_{110}(1-Q_{11})-P_{010}(1-Q_{01})}{1-Q_{*}}$
$\qquad+\frac{\min\left\{0,Q_{*}-P_{111}Q_{11}\right\}+\min\left\{P_{011}Q_{01},Q_{01}-Q_{*}\right\}}{Q_{*}(1-Q_{*})}.$
We provide the derivation of these bounds (and those in the next proposition) in Section A.3. These bounds will differ from the above randomization bounds given in Proposition 3 in two ways. First, with the randomization assumption alone, we could only leverage the observed strata within levels of treatment—further stratification in terms of the moderator provided no information because it placed no restriction on the relationship between the pre-test and post-test versions of the moderator. Under moderator monotonicity (Assumption 4), we can leverage $P_{tzd}$ and $Q_{tz}$ to narrow the bounds. Second, moderator monotonicity places bounds on the true value of $Q_{*}$ since it must be less than $\min(Q_{11},Q_{01})$.
While the moderator monotonicity assumption does narrow the bounds, they are often still quite wide and usually contain 0. To further narrow the bounds, we consider another assumption that the moderator is stable in the control arm of the study.
Assumption 5 (Stable Moderator under Control).
$M^{*}_{i}=M_{i}(0)$
This assumption implies that the moderator under control in the post-test design is the same as the moderator as if it was measured pre-test. This assumption may be plausible in experimental designs where the control condition is neutral or similar to the pre-test environment. In our empirical example, this would mean that hearing the generic campaign speech, which does not mention the land issue, does not affect perceived land insecurity. Under both Assumptions 4 and 5, the only values that principal strata that $S_{i}$ can take are $\{111,100,000\}$.
Proposition 5 (Post-test Sharp Bounds under Moderator Monotonicity and Stability).
Under Assumptions 3, 4, and 5, we have $Q_{*}=Q_{01}$ and sharp bounds $\delta\in[\delta_{L3},\delta_{U3}]$, where
$\delta_{L3}$ $=\frac{P_{111}Q_{11}}{Q_{01}}-P_{011}-\frac{P_{110}(1-Q_{11})}{1-Q_{01}}+P_{010}-\min\left\{1,\frac{P_{111}Q_{11}}{Q_{11}-Q_{01}}\right\}\cdot\frac{Q_{11}-Q_{01}}{Q_{01}(1-Q_{01})},$
$\delta_{U3}$ $=\frac{P_{111}Q_{11}}{Q_{01}}-P_{011}-\frac{P_{110}(1-Q_{11})}{1-Q_{01}}+P_{010}-\max\left\{0,\frac{P_{111}Q_{11}-Q_{01}}{Q_{11}-Q_{01}}\right\}\cdot\frac{Q_{11}-Q_{01}}{Q_{01}(1-Q_{01})}.$
These bounds demonstrate how the magnitude of treatment effect on the moderator affects identification in the post-test design. A unit’s moderator is affected by treatment whenever $M_{i}(1)\neq M_{i}(0)$, which corresponds to the $S_{i}=100$ principal stratum under monotonicity and stable moderator under control. Note that under these assumptions, $\rho_{100}$ represents the magnitude of the treatment-moderator effect. Since $Q_{11}=\rho_{111}+\rho_{100}$ and $Q_{01}=\rho_{111}$, we can identify this effect with the usual difference in (population) means, $Q_{11}-Q_{01}=\rho_{100}$. The maximum possible width of the sharp bounds depends on this effect, with
$\max\{\delta_{U3}-\delta_{L3}\}=\frac{Q_{11}-Q_{01}}{Q_{01}(1-Q_{01})}=\frac{\rho_{100}}{\rho_{000}(\rho_{111}+\rho_{100})},$
so that the bounds can be relatively informative if the post-treatment average effect on the moderator is small.
4.2.2 Sensitivity Analysis under Limited Effects on the Moderator
While the monotonicity and stable moderator assumptions can considerably narrow the nonparametric bounds on our causal quantities, they rule out entire principal strata, which may be stronger than is justified for a particular empirical setting. We now consider an alternative approach to bounds that does not rule out any particular principal strata but rather places restrictions on the proportion of units whose moderators are affected by treatment.
In particular, we propose a sensitivity analysis that limits the proportion of respondents whose moderator value changes between the pre-test and post-test, regardless of the treatment condition (contrast this with Assumption 5 which applies to the control condition only). We operationalize this via the following constraint,
$\Pr(S_{i}\notin\{111,000\})\leq\gamma.$
Note that $\gamma$ must be greater than $|Q_{11}-Q_{01}|$ for the bounds to be feasible since $|Q_{11}-Q_{01}|=|\rho_{101}+\rho_{100}-\rho_{011}-\rho_{010}|$. We vary the value of $\gamma$ from $|Q_{11}-Q_{01}|$ to $1$ and see how the nonparametric bounds on the value of $\delta$ change as we gradually allow a larger treatment effect on the moderator. We obtain these new bounds by adding this additional constraint to linear program, which we can then solve numerically. This approach is a more flexible way to allow for limited heterogeneous treatment effects on the moderator in any direction. Researchers can also combine this sensitivity analysis with the monotonicity and stable moderator assumptions.
4.3 Randomized Placement Design
Finally, we examine a combined pre/post design called the randomized placement design, where in addition to treatment, the timing of moderator measurement, $Z_{i}$, is also randomized. Bounds from this design will improve upon the post-test bounds because we can identify the marginal probability of the true moderator as
$Q_{*}\ =\ \Pr(M^{*}_{i}=1)\ =\ \Pr(M_{i}=1\mid Z_{i}=0).$
Unfortunately, for the randomization-only bounds in Equation (12), the sign of $\delta$ cannot be identified for any value of $Q_{*}\in(0,1)$.
With the randomized placement design, we have a slightly more complicated set of principal strata since now we must handle both the pre-test and post-test potential outcomes. In particular, we have the following:
$\psi_{y_{1}y_{0}s}(t)\ =\ \Pr(Y^{*}_{i}(t)=y_{1},Y_{i}(t)=y_{0}\mid S_{i}=s)$
where $\psi_{y_{1}y_{0}s}(t)\geq 0$ and $\sum_{y_{1}}\sum_{y_{0}}\psi_{y_{1}y_{0}s}(t)=1$ for all $t$. These values characterize the joint distribution of the pre-test and post-test potential outcomes for a given treatment level, $t$, and principal strata, $s$. Furthermore, let $\mathcal{S}^{*}_{m}=\{s:s\in\mathcal{S},M^{*}_{i}=m\}$ be the set of principal strata with the true value of the moderator equal to $m$.
Given these, we can write the interaction between the treatment and the moderator as
$\delta\ =\ \sum_{y_{0}=0}^{1}\left\{\sum_{s_{1}\in\mathcal{S}^{*}_{1}}\frac{\rho_{s_{1}}}{Q_{*}}(\psi_{1y_{0}s_{1}}(1)-\psi_{1y_{0}s_{1}}(0))-\sum_{s_{0}\in\mathcal{S}^{*}_{0}}\frac{\rho_{s_{0}}}{1-Q_{*}}(\psi_{1y_{0}s_{0}}(1)-\psi_{1y_{0}s_{0}}(0))\right\},$
and the observed strata in the pre-test and post-test arms as
$P_{t1m}Q_{t1}$ $=\Pr(Y_{i}=1,D_{i}=m\mid T_{i}=t,Z_{i}=1)\ =\ \sum_{y_{0}=0}^{1}\sum_{s\in\mathcal{S}(t,1,m)}\psi_{1y_{0}s}(t)\rho_{s},$
$P_{t0m_{\ast}}$ $=\Pr(Y_{i}=1\mid T_{i}=t,Z_{i}=0,M_{i}=m_{\ast})=\frac{\sum_{y_{1}=0}^{1}\sum_{s\in\mathcal{S}^{*}_{m_{\ast}}}\psi_{y_{1}1s}(t)\rho_{s}}{\sum_{s\in\mathcal{S}^{*}_{m_{\ast}}}\rho_{s}},$
for all values of $y$, $m$, $t$, and $m_{\ast}$.
Without further assumptions, the pre-test data are helpful only insofar as they identify the marginal distribution of the moderator, $Q_{*}$. We can narrow the bounds under the random placement design using all of the substantive assumptions described above: priming monotoncity, moderator monotonicity, and stability under control. Each of these implies restrictions on the above principal strata that can be incorporated into a linear programming problem that can be solved numerically to provide the bounds.
4.3.1 Sensitivity Analysis in the Random Placement Design
The random placement design allows us to combine the sensitivity analysis procedure of the pre- and post-test designs. In particular, we can impose both of the restrictions from above simultaneously:
$\Pr(S_{i}\notin\{111,000\})$ $\leq\ \gamma,$
$\Pr(Y^{*}_{i}(t)=1,Y_{i}(t)=0\mid M^{*}_{i}=m_{\ast})+\Pr(Y^{*}_{i}(t)=0,Y_{i}(t)=1\mid M^{*}_{i}=m_{\ast})$ $\leq\ \theta$
Again, these restrict (a) how much the treatment and moderator measurement timing affect the moderator, and (b) how much the moderator measurement timing affects the outcome. To incoprate these restrictions into the bounds, we can rewrite them in the principal strata described above:
$(1-\rho_{111}-\rho_{000})$ $\leq\ \gamma,$
$\frac{\sum_{s\in\mathcal{S}^{*}_{m_{\ast}}}\left(\psi_{10s}(t)+\psi_{01s}(t)\right)\rho_{s}}{\sum_{s\in\mathcal{S}^{*}_{m_{\ast}}}\rho_{s}}$ $\leq\ \theta.$
We can easily add these restrictions to the optimization problem that produces the bounds on the interaction.
We can conduct sensitivity analyses on the randomized placement design by varying the values of $\gamma$ and $\theta$ and seeing how the values of the bounds change. There are several ways to conduct and present such a two-dimensional sensitivity analysis. One would be to plot the parameters on each axis and demarcate the regions where the bounds are informative (e.g., do not include zero) and where they are not. A second approach would be to choose a small value for one of the two parameters consistent with a researcher’s beliefs. For instance, if a researcher believes that the moderator is unlikely to be affected by treatment, then they could choose a small value for $\gamma$ and investigate the sensitivity of the bounds to different amounts of priming, as measured by $\theta$.
4.4 Statistical Inference
The above bounds are stated in terms of population quantities when, in fact, we only ever have sample data. We can easily obtain estimates for the bounds by plugging in sample versions of these probabilities or, for the random placement design, solving the linear programming problem using the sample data. Obtaining valid confidence intervals in this setting is more challenging, since standard asymptotic analyses break down due to the maximum and minimum operators causing nondifferentiability. This problem can lead the standard nonparametric bootstrap to have problematic theoretical properties (Andrews and Han, 2009; Fang and Santos, 2018).
Furthermore, confidence intervals for the bounds tend to be overly conservative when the target of inference is the parameter $\delta$ rather than the bounds themselves. As pointed out by Imbens and Manski (2004), this occurs because the true parameter cannot simultaneously be close to the upper and lower bound at the same time. This fact allows us to narrow the confidence intervals by a data-driven amount while maintaining nominal coverage for the parameter of interest.
Our approach to inference combines the Imbens and Manski (2004) approach with estimated standard errors of the bounds from the the nonparametric bootstrap. While the bootstrap may have theoretical problems, we find in simulations that this approach produces conservative confidence intervals that have slightly higher-than-nominal empirical coverage. In Supplemental Materials B, we provide further details of our approach to estimation. In Supplemental Materials D, we also develop a Bayesian model-based approach to incorporate additional pre-treatment covariates that might be available to researchers.
5 Empirical Example
To illustrate how our approach can be applied to both the post-test and randomized placement design, we return to the example from Horowitz and Klaus (2020) introduced in Section 2. Recall that the interaction in this case shows how the effect of land-based appeals varies by a respondent’s level of land security.
5.1 Setup and Assumptions
First, it is worth discussing the assumptions beyond randomization in this context. Assumptions 4 (moderator monotonicity) and 5 (stability) are not guaranteed by the design and require a substantive justification. In this case, moderator monotonicity requires the placement of the land insecurity question after treatment shift perceived land insecurity in the same direction for all respondents (or have no effect). We assume a positive (or zero) individual effect on land insecurity, consistent with the estimate ATE—though it is substantively small and not statistically significant ($\hat{\beta}=0.04$, $p=0.23$). Given that one part of the active treatment was intended to induce fear of “land grabbing,” it seems somewhat plausible that measuring perceived land rights after treatment would only increased feelings of insecurity.
The assumption of stability means that hearing the generic campaign speech (with no appeals to the land issue) has no individual-level effect on perceived land insecurity when it is measured post-treatment. Again, this assumption cannot be conclusively tested with the observed data since we cannot estimate individual-level effects. However, with the randomized placement design, we can estimate the ATE of the pre/post randomization on the moderator among respondents in the control group. Our estimate of the ATE is very small and not statistically significant ($\hat{\beta}=-0.005$, $p=0.91$), which is at least consistent with the assumption of a stable moderator under control.
5.2 Comparing the Sharp Bounds across Different Assumptions
Figure 2: Estimated nonparametric bounds (thick bars) and 95% confidence intervals (thin bars) under different designs and assumptions. The final panel contains OLS point estimates and 95% confidence intervals.
Figure 2 displays the non-parametric bounds (with 95% confidence intervals) for $\delta$ under different sets of assumptions applied to the post-test data (in grey), the pre-test (in blue), and the combined random placement design (labeled “Prepost” in black). We also include the naïve OLS estimate with 95% confidence interval for comparison. Some of the designs are missing from panels because an assumption does not apply to that design (for example, moderator monotonicity under the pre-test design).
Assuming only randomization, the nonparametric bounds are uninformative of the sign of $\delta$, and are much wider than the confidence interval of the naïve OLS estimate for all of the designs. Adding the assumption of moderator monotonicity reduces the width of the bounds, especially on the combined prepost data. Unfortunately, the confidence intervals become considerably larger, especially for the post-test data, in which the $M_{i}=1$ group is small.
The assumption of stability tightens the bounds to a similar degree for the post-only and pre-post data. Under the pre-test assumptions of randomization, moderator monotonicity, and stability, the bounds exclude zero for both the post-test and random placement designs, though the confidence interval for the latter contains zero. Furthermore, the bounds for this design are wider under all three assumptions than under just randomization and monotonicity since, in finite samples, the pre-test mean of $M_{i}$ differs from the post-test mean of $M_{i}$ in the control arm. We would expect a difference by random chance even when stability holds, but the bounds may widen slightly to accommodate this divergence between the population and sample quantities.
Priming monotonicity also narrows the bounds for the pre-test and random placement designs, though not by as much as moderator monotonicity. The combination of the two monotonicity assumptions, however, recovers bounds that are close to the OLS estimate and has confidence intervals that barely contain zero.
In sum, the nonparametric bounds do not support the hypothesis of a positive interaction effect under the randomization assumptions alone. It is only with a combination of additional substantive assumptions—moderator monotonicity, priming monotonicity, and stability—that the sharp bounds produce results qualitatively similar to the naïve OLS estimate of a positive interaction effect, though the designs differ on whether this is statistically significant or not.
5.3 Implementing the Sensitivity Analysis
We now apply our sensitivity analysis procedures to this experiment. These procedures involve varying the proportion of respondents for whom the placement of the moderator measure affects their moderator value (land security) or their outcome (candidate support), labeled $\gamma$ and $\theta$ respectively. We can apply the $\gamma$ and $\theta$ sensitivity analyses to the post-test and pre-test designs, respectively, and we can combine them in the random placement design.
Figure 3 shows the post-test bounds as a function of $\gamma$ under just the randomization assumption. While $\gamma$ can theoretically range up to 1, here we limit it to 0.5 to aid presentation since the bounds quickly stabilize. The black lines denote the upper and lower bounds, and the shaded ribbon denotes the 95% confidence intervals around the bounds. The lower bound crosses 0 when $\gamma=0.07$—that is, when no more than 7% of respondents are affected by the post-treatment measurement of the moderator. The 95% confidence interval contains 0 even for the minimum possible value of $\gamma$ consistent with the observed data (0.02). The sensitivity analysis shows that the sharp bounds are highly sensitive to changes in the degree of post-treatment bias for small values of $\gamma$.
Figure 3: Post-test sensitivity analysis. Nonparametric bounds (black lines) with 95% confidence intervals (grey ribbon) as a function of $\gamma$, the proportion of respondents whose value of the moderator variable (land insecurity) is affected by post-test measurement.
To interpret this sensitive analysis, researchers will need to draw on their substantive knowledge to assess the plausible range of $\gamma$. The estimated ATE on the post-treatment moderator is $0.02$ ($p=0.59$), but without a monotonicity assumption this may include respondents with positive and negative effects that offset their effects. We might assume that most of the effect of the land rights prime would accrue to respondents with less certain responses such as feeling “somewhat secure” in their land rights and that had been personally affected by prior ethnic conflict. Using the pre-test data, we find that this is 12% of the sample, which we might consider a reasonable value of $\gamma$. While it may seem like a relatively minor problem if only 12% of the sample is affected, our sensitivity analysis shows that the nonparametric bounds would be about eight times wider ($[-1.02,1.14]$) compared to the case where $\gamma$ is at its minimum ($[0.38,0.64]$).
Figure 4: Pre-test sensitivity analysis. Nonparametric bounds (black lines) with 95% confidence intervals (grey ribbon) as a function of $\theta$, the proportion of respondents who are primed by asking the moderator before treatment, under priming monotoncity assumption.
For priming bias in the pre-test design, we can assess sensitivity as a function of $\theta$, the proportion of respondents whose outcome value is affected by when the moderator is measured. Figure 4 shows the bounds as a function of $\theta$ under priming monotonicity. When $\theta=0$, we are assuming that priming bias does not exist so the true interaction is point identified in the pre-test arm, though the confidence interval at $\theta=0$ already includes zero. The estimated lower bound crosses zero at approximately 0.2, when up to 20% of respondents are primed, which would be a rather large effect of priming.
Finally, in the random placement design, we combine these two tests by restricting both $\gamma$ and $\theta$. Figure 5 shows the prepost bounds as a function of $\gamma$ under two different assumption about the amount of priming: limited priming $\theta\leq 0.25$ and unrestricted priming $\theta\leq 1$. In both settings, the bounds initially shrink as $\gamma$ increases, most likely to due to the low values of $\gamma$ being inconsistent with the data. The bounds then begin to widen, though they widen much more for the unrestricted priming compared the limited priming setting.
Figure 5: Random placement design sensitivity analysis. Nonparametric bounds (black lines) and 95% confidence intervals (grey ribbons) as a function of the post-test effect on the moderators and the amount of priming. The limited priming assumption assumes $\theta\leq 0.25$ and the unrestricted priming has $\theta\leq 1$.
6 Concluding Remarks
This paper addresses a central tension in survey methodology: how should researchers assess priming bias versus post-treatment bias when designing a survey experiment? We provide sharp bounds for interactions with moderators measured post-treatment and show how these bounds vary under additional substantive assumptions. We also provide sensitivity analyses for both types of bias by varying the proportion of respondents whose moderator value changes in the post-test design and the proportion of respondents for whom the pre-test measurement of the moderator would prime their responses. We demonstrate how these tools can be used to diagnose and assess the severity of post-treatment bias and priming bias by applying them to a survey experiment regarding the effect of land-based appeals by politicians on electoral support in Kenya.
Open questions remain from our approach here. In particular, future work could optimize the random placement design to balance the priming and post-treatment bias concerns. In addition, we could consider how integrating separate pre-test surveys, often given weeks or months before treatment, might allow for a different set of plausible assumptions and identification.
Appendix A Proofs
A.1 Pre-test bounds
Proof of Proposition 4.
Note that we can write $\tau_{pre}(m)=P_{10m}-P_{00m}$. From expression 7, we can write the true CATE under priming monotonicity as
$\tau(m)$ $=P_{10m}-P_{00m}-\left\{\mathbb{P}(Y^{*}_{i}(1)=0,Y_{i}(1)=1\mid M^{*}_{i}=m)-\mathbb{P}(Y^{*}_{i}(0)=0,Y_{i}(0)=1\mid M^{*}_{i}=m)\right\}.$
We can bound the unknown probabilities as
$0\leq\mathbb{P}(Y^{*}_{i}(t)=0,Y_{i}(t)=1\mid M^{*}_{i}=m)\leq P_{t0m}.$
Thus, a sharp upper bound on $\tau(m)$ would because
$P_{10m}-P_{00m}+P_{00m}=P_{10m},$
and a sharp lower bound would be
$P_{10m}-P_{00m}-P_{10m}=-P_{00m},$
which establishes the bounds for $\tau(m)$. For the upper $\delta=\tau(1)-\tau(0)$, we simply use the upper bound for $\tau(1)$ and the lower bound for $\tau(0)$ and vice versa for the lower bound for $\delta$. ∎
Proof of Proposition 2.
Under priming mononoticity, we can write the true CATE as
$\tau(m)$ $=\tau_{pre}(m)-\left\{\mathbb{P}(Y^{*}_{i}(1)=0,Y_{i}(1)=1\mid M^{*}_{i}=m)-\mathbb{P}(Y^{*}_{i}(0)=0,Y_{i}(0)=1\mid M^{*}_{i}=m)\right\}.$
The restriction on the amount of priming bias implies that
$0\leq\mathbb{P}(Y^{*}_{i}(t)=0,Y_{i}(t)=1\mid M^{*}_{i}=m)\leq\theta,$
for all $t$. Thus, clearly we have
$\tau(m)\in\left[\tau_{pre}(m)-\theta,\tau_{pre}(m)+m\right].$
Taking the maximum and minimum of $\tau(1)$ and $\tau(0)$, respectively, establishes the upper bound for $\delta$. Reversing this gives the lower bound.
A.2 Post-test randomization bounds
Proof of Proposition 3.
Under Assumption 3, the information about the parameter of interest comes from $P_{t}=\mathbb{P}(Y_{i}=1\mid T_{i}=t,Z_{i}=1)$ alone. This is because the distribution of the post-test moderators provides no information about the pre-test moderator. Recall that
$P_{t}=\pi_{t1}Q_{*}+\pi_{t0}(1-Q_{*}),$ (13)
where $\pi_{tm}=\mathbb{P}[Y_{i}(t,1)=1\mid M^{*}_{i}=m]$ and $Q_{*}=\mathbb{P}[M^{*}_{i}=1]$.
Below, we show how to derive the upper bound for $\delta$. The derivation of the lower bound is similar. Conditional on $Q_{*}$, we can define the following linear program:
$\max$ $\pi_{11}-\pi_{01}-\pi_{10}+\pi_{00}$
subject to $\pi_{t1}Q_{*}+\pi_{t0}(1-Q_{*})=P_{t}\quad\text{for }t=0,1,$
$0\leq\pi_{tm}\leq 1\quad\forall(t,d)\in\{0,1\}^{2}$
We can convert this to an augmented form by adding slack variables,
$\max$ $\pi_{11}-\pi_{01}-\pi_{10}+\pi_{00}$
subject to $\pi_{t1}Q_{*}+\pi_{t0}(1-Q_{*})=P_{t}\quad\text{for }t=0,1$
$\pi_{tm}+s_{tm}=1\quad\forall(t,m)\in\{0,1\}^{2}$
$\{\pi_{11},\pi_{01},\pi_{10},\pi_{00},s_{11},s_{01},s_{10},s_{00}\}\geq 0.$
The feasibility of various basic solutions here will depend on the relationship between the observed probabilities and $Q_{*}$. In Table SM.1, we show basic feasible solutions for the four different conditions relating $P_{1}$ and $P_{0}$ to $Q_{*}$. Under each condition, it is straightforward to determine that the basic feasible solution is also optimal since there is no entering variable that can increase the value of the quantity of interest. Thus, we know that
$\delta\leq\min\left\{\frac{1+P_{1}-P_{0}}{Q_{*}},\ \frac{1-P_{1}+P_{0}}{1-Q_{*}},\ \frac{1-P_{0}}{Q_{*}}+\frac{1-P_{1}}{1-Q_{*}},\ \frac{P_{1}}{Q_{*}}+\frac{P_{0}}{1-Q_{*}}\right\}$
Table SM.1: Optimal solutions to the linear program under different conditions
Condition $\pi_{11}$ $\pi_{10}$ $\pi_{01}$ $\pi_{00}$ $s_{11}$ $s_{01}$ $s_{10}$ $s_{00}$ Value
$P_{1}>Q_{*},P_{0}>1-Q_{*}$ $\frac{P_{1}}{Q_{*}}$ 0 0 $\frac{P_{0}}{1-Q_{*}}$ $1-\frac{P_{1}}{Q_{*}}$ 1 1 $1-\frac{P_{0}}{1-Q_{*}}$ $\frac{P_{1}}{Q_{*}}+\frac{P_{0}}{1-Q_{*}}$
$P_{1}<Q_{*},P_{0}>1-Q_{*}$ $1$ $\frac{P_{1}-Q_{*}}{1-Q_{*}}$ 0 $\frac{P_{0}}{1-Q_{*}}$ 00 1 1 $1-\frac{P_{0}}{1-Q_{*}}$ $\frac{1-P_{1}+P_{0}}{1-Q_{*}}$
$P_{1}>Q_{*},P_{0}<1-Q_{*}$ $\frac{P_{1}}{Q_{*}}$ 0 $\frac{P_{0}-1+Q_{*}}{Q_{*}}$ $1$ $1-\frac{P_{1}}{Q_{*}}$ 1 1 00 $\frac{P_{1}-P_{0}+1}{Q_{*}}$
$P_{1}<Q_{*},P_{0}<1-Q_{*}$ $1$ $\frac{P_{1}-Q_{*}}{1-Q_{*}}$ $\frac{P_{0}-1+Q_{*}}{Q_{*}}$ $1$ 00 1 1 00 $\frac{1-P_{1}}{1-Q_{*}}+\frac{1-P_{0}}{Q_{*}}$
A similar derivation shows that
$\delta\geq\max\left(-\frac{1-P_{1}+P_{0}}{Q_{*}},\ -\frac{1+P_{1}-P_{0}}{1-Q_{*}},\ -\frac{P_{0}}{Q_{*}}-\frac{P_{1}}{1-Q_{*}},\ -\frac{1-P_{1}}{Q_{*}}-\frac{1-P_{0}}{1-Q_{*}}\right).$
Taking a look at these bounds, suppose that $P_{1}\geq 1-P_{0}$, then the upper bound as a function of $Q_{*}$ is:
$U(Q_{*})=\begin{cases}\frac{P_{0}+(1-P_{1})}{1-Q_{*}}&\text{if}\;Q_{*}\leq 1-P_{0}\\ \frac{1-P_{0}}{Q_{*}}+\frac{1-P_{1}}{1-Q_{*}}&\text{if}\;Q_{*}\in[1-P_{0},P_{1}]\\ \frac{P_{1}+(1-P_{0})}{Q_{*}}&\text{if}\;Q_{*}\geq P_{1}\end{cases}$
Notice that the numerators of all these functions are positive, so the first bounding function $(P_{0}+1-P_{1})/(1-Q_{*})$ is montonically increasing over its range and the third, $(P_{1}+(1-P_{0}))/Q_{*}$ is monotonically decreasing over its range. Finally, inspection of the middle bounding function shows that it is convex over its range. This implies that the this function must have its maximum at one of the bound points, $1-P_{0}$ or $P_{1}$. Taking the maximum of these two values, and comparing them to the maximum two values from the situation when $P_{1}\leq 1-P_{0}$ gives the expression of the bounds in the result. A similar approach applies to the lower bounds as well.
Sharpness of these bounds is implied by the linear nature of the optimization function and the convexity of the feasible set. If these bounds were not sharp, this would imply that that there are bounds sharper than these that contain all values of $\delta$ consistent with the data and maintained assumptions. But this is clearly contradicted by the fact that the solutions in Table SM.1 are feasible and would fall outside these supposedly sharper bounds. ∎
A.3 Bounds under additional assumptions
To derive bounds under additional assumptions, we first derive bounds conditional on the strata probabilities.
Lemma SM.1.
The bounds on $\delta$ for given values of $\bm{\rho}$ are $\delta\in[\delta_{L}(\bm{\rho}),\delta_{U}(\bm{\rho})]$, where
$\delta_{U}(\bm{\rho})$ $=P_{111}\left(\frac{Q_{11}}{Q_{*}}\right)-P_{011}\left(\frac{Q_{01}}{Q_{*}}\right)-P_{110}\left(\frac{1-Q_{11}}{1-Q_{*}}\right)+P_{010}\left(\frac{1-Q_{01}}{1-Q_{*}}\right)$
$+\frac{1}{Q_{*}(1-Q_{*})}\min\left\{\begin{matrix}P_{110}(1-Q_{11})\\ \rho_{011}+\rho_{001}\\ Q_{*}-P_{111}Q_{11}\end{matrix}\right\}+\frac{1}{Q_{*}(1-Q_{*})}\min\left\{\begin{matrix}P_{011}Q_{01}\\ \rho_{110}+\rho_{010}\\ 1-Q_{*}-P_{010}(1-Q_{01})\end{matrix}\right\},$
and,
$\delta_{L}(\bm{\rho})$ $=P_{111}\left(\frac{Q_{11}}{Q_{*}}\right)-P_{011}\left(\frac{Q_{01}}{Q_{*}}\right)-P_{110}\left(\frac{1-Q_{11}}{1-Q_{*}}\right)+P_{010}\left(\frac{1-Q_{01}}{1-Q_{*}}\right)$
$+\frac{1}{Q_{*}(1-Q_{*})}\max\left\{\begin{matrix}-P_{111}Q_{11}\\ -\rho_{110}-\rho_{100}\\ -1+Q_{*}+P_{110}(1-Q_{11})\end{matrix}\right\}+\frac{1}{Q_{*}(1-Q_{*})}\max\left\{\begin{matrix}-P_{011}Q_{01}\\ -\rho_{001}-\rho_{101}\\ -Q_{*}+P_{011}Q_{01}\end{matrix}\right\}.$
Proof.
Conditional on $\rho_{s}$ and $Q_{*}$, deriving the bounds on $\delta$ is a standard linear programming problem. We now describe the process for deriving these bounds at a general level. Without any assumptions, we are interested in maximizing or minimizing the objective function,
$\delta=$ $P_{111}\left(\frac{Q_{11}}{Q_{*}}\right)-P_{011}\left(\frac{Q_{01}}{Q_{*}}\right)-P_{110}\left(\frac{1-Q_{11}}{1-Q_{*}}\right)+P_{010}\left(\frac{1-Q_{01}}{1-Q_{*}}\right)$
$+\mu_{011}(1,1)\frac{\rho_{011}}{Q_{*}(1-Q_{*})}+\mu_{001}(1,1)\frac{\rho_{001}}{Q_{*}(1-Q_{*})}$
$-\mu_{110}(1,1)\frac{\rho_{110}}{Q_{*}(1-Q_{*})}-\mu_{100}(1,1)\frac{\rho_{100}}{Q_{*}(1-Q_{*})}$
$+\mu_{110}(0,1)\frac{\rho_{110}}{Q_{*}(1-Q_{*})}+\mu_{010}(0,1)\frac{\rho_{010}}{Q_{*}(1-Q_{*})}$
$-\mu_{101}(0,1)\frac{\rho_{101}}{Q_{*}(1-Q_{*})}-\mu_{001}(0,1)\frac{\rho_{001}}{Q_{*}(1-Q_{*})},$
subject to the constraints
$P_{111}Q_{11}$ $=\mu_{111}(1,1)\rho_{111}+\mu_{101}(1,1)\rho_{101}+\mu_{110}(1,1)\rho_{110}+\mu_{100}(1,1)\rho_{100}$
$P_{011}Q_{01}$ $=\mu_{111}(0,1)\rho_{111}+\mu_{011}(0,1)\rho_{011}+\mu_{110}(0,1)\rho_{110}+\mu_{010}(0,1)\rho_{010}$
$P_{110}(1-Q_{11})$ $=\mu_{011}(1,1)\rho_{011}+\mu_{110}(1,1)\rho_{110}+\mu_{010}(1,1)\rho_{010}+\mu_{000}(1,1)\rho_{000}$
$P_{010}(1-Q_{01})$ $=\mu_{101}(0,1)\rho_{101}+\mu_{001}(0,1)\rho_{001}+\mu_{100}(0,1)\rho_{100}+\mu_{000}(0,1)\rho_{000}$
00 0 $\leq\mu_{s}(t)\leq 1,\quad\forall s,t.$
For this step, we do not need to specify constraints on $\rho_{s}$ because we consider them fixed (and $Q_{*}$ is a linear function of $\rho_{s}$). The simplex tableau method yields the given bounds. ∎
Proof of Proposition 4.
Recall the constraints on the strata probabilities:
$Q_{11}$ $=\rho_{111}+\rho_{101}+\rho_{110}+\rho_{100}$
$Q_{01}$ $=\rho_{111}+\rho_{011}+\rho_{110}+\rho_{010}$
$Q_{*}$ $=\rho_{111}+\rho_{011}+\rho_{101}+\rho_{001},$
Under monotonicity, we only have strata $S_{i}\in\{111,110,010,100,000\}$, so we have $Q_{*}=\mathbb{P}[M^{*}_{i}=1]=\rho_{111}$ and $\rho_{110}+\rho_{010}=Q_{01}-Q_{*}$ and $\rho_{110}+\rho_{100}=Q_{11}-Q_{*}$. Plugging these values into the bounds from Lemma SM.1, we obtain
$\delta_{U}(Q_{*})$ $=P_{111}\left(\frac{Q_{11}}{Q_{*}}\right)-P_{011}\left(\frac{Q_{01}}{Q_{*}}\right)-P_{110}\left(\frac{1-Q_{11}}{1-Q_{*}}\right)+P_{010}\left(\frac{1-Q_{01}}{1-Q_{*}}\right)$
$+\frac{1}{Q_{*}(1-Q_{*})}\min\left\{\begin{matrix}P_{110}(1-Q_{11})\\ 0\\ Q_{*}-P_{111}Q_{11}\end{matrix}\right\}+\frac{1}{Q_{*}(1-Q_{*})}\min\left\{\begin{matrix}P_{011}Q_{01}\\ Q_{01}-Q_{*}\\ 1-Q_{*}-P_{010}(1-Q_{01})\end{matrix}\right\},$
and,
$\delta_{L}(Q_{*})$ $=P_{111}\left(\frac{Q_{11}}{Q_{*}}\right)-P_{011}\left(\frac{Q_{01}}{Q_{*}}\right)-P_{110}\left(\frac{1-Q_{11}}{1-Q_{*}}\right)+P_{010}\left(\frac{1-Q_{01}}{1-Q_{*}}\right)$
$+\frac{1}{Q_{*}(1-Q_{*})}\max\left\{\begin{matrix}-P_{111}Q_{11}\\ -Q_{11}-Q_{*}\\ -1+Q_{*}+P_{110}(1-Q_{11})\end{matrix}\right\}+\frac{1}{Q_{*}(1-Q_{*})}\max\left\{\begin{matrix}-P_{011}Q_{01}\\ 0\\ -Q_{*}+P_{011}Q_{01}\end{matrix}\right\}.$
We further simplify the upper bound expression by noting that $P_{110}(1-Q_{11})\geq 0$ and $Q_{01}-Q_{*}\leq 1-Q_{*}-P_{010}(1-Q_{01})$. The lower bound simplifies because $Q_{11}-Q_{*}\leq 1-Q_{*}-P_{110}(1-Q_{11})$ and $P_{011}Q_{01}\geq 0$. Removing these extraenous conditions gives the result in the text.
Proof of Proposition 5.
Under the maintained assumptions, $Q_{*}=Q_{01}$, which we plug into the expression of Proposition 4. Then, the result is immediate upon noting that $P_{011}Q_{01}-Q_{01}\leq 0$, $P_{011}Q_{01}\geq 0$ and rearranging terms. ∎
For calculating the bounds under the sensitivity constraints, we can take the bounds from Lemma SM.1 and solve a corresponding linear programming problem to optimize them with respect to the principal strata probabilities. For example, depending the observed data, the upper bound will depend on $\rho_{011}+\rho_{001}$, $\rho_{110}+\rho_{010}$, or $\rho_{011}+\rho_{001}+\rho_{110}+\rho_{010}$. To find the upper bound across values of $\bm{\rho}$, we apply the linear programming machinery to finding the upper bound for each of these quantities subject to the constraints that
$Q_{11}$ $=\rho_{111}+\rho_{101}+\rho_{110}+\rho_{100}$
$Q_{01}$ $=\rho_{111}+\rho_{011}+\rho_{110}+\rho_{010}$
$Q_{*}$ $=\rho_{111}+\rho_{011}+\rho_{101}+\rho_{001},$
where $0\leq\rho_{s}\leq 1$ for all $s$ and $\sum_{s\in\mathcal{S}}\rho_{s}=1$. Note that for the sensitivity analysis, we may impose additional constraints on $\rho_{s}$ in this step. As an example, for the objection function of $\rho_{011}+\rho_{001}$, we have the upper bound
$\min\left\{1-Q_{11},Q_{*},1-Q_{01}+Q_{*},\frac{1}{2}(1-Q_{11}+Q_{*}),1+Q_{01}-Q_{11}\right\}.$
Plugging these bounds into the upper bound $\delta_{U}(\bm{\rho})$ will yield an upper bound purely as a function of observed parameters and $Q_{*}$ and $\gamma$ ( the sensitivity parameter). Under some of our assumptions, inspection of the resulting functions reveals that the maximum of these functions can only occur at a handful of critical values of $Q_{*}$ which can be evaluated and compared quickly. Otherwise, we use a standard optimization routine to find the value of $Q_{*}$ that maximizes the upper bound or minimizes the lower bound.
Appendix B Estimation and Inference Details
The discussion of the bounds in the main text focused on population level bounds—that is, we identified the bounds in terms of population quantities such as $P_{tzm}$. Estimation and inference for the bounds with a sample poses some important difficulties. The most obvious way to estimate these bounds is to plug in sample version of the population quantities and solve the above linear programming problem to obtain bounds. Unfortunately, the asymptotic distribution of estimators based on this plug-in approach do not have the standard asymptotics due to the lack of differentiability of the bounds as a function of the data. Andrews and Han (2009) show that naive bootstrap methods are not valid for these types of estimators due to these issues.
Below we mostly focus on the random placement design, where estimation and inference is the most complicated. In the pre-test design, the bounds are simple functions of the parameters of interest, so we can use standard asymptotic variance estimators, combined with the approach of Imbens and Manski (2004) to obtain confidence intervals for the parameters of interest. For the post-test bounds, we obtain standard errors for the bounds based on the nonparametric bootstrap and then use these in the approach of Imbens and Manski (2004). In simulations, we found this to have similar performance to the more complicated union bound approach of Ye et al. (2023).
B.1 Estimation and inference for the random placement design
For the random placement design, we do not generally have closed-form expressions for the bounds, and so we can reformulate the estimation and inference problem based on moment conditions that feed into sample versions of a population criterion function (Chernozhukov, Hong and Tamer, 2007). In the general analysis of the random placement design, we define the parameters of our model as
$\psi_{y_{1}y_{0},m_{1}}(t,m_{0})=\Pr[Y^{*}_{i}(t)=y_{1},Y_{i}(t)=y_{0},M_{i}(t)=m_{1}\mid M_{i}^{*}=m_{0}],$
so that the parameter of interest can be written as
$\delta=\sum_{y_{0}m_{1}}\psi_{1y_{0}m_{1}}(1,1)-\psi_{1y_{0}m_{1}}(0,1)-\psi_{1y_{0}m_{1}}(1,0)+\psi_{1y_{0}m_{1}}(0,0),$
with constraints
00 0 $\leq\psi_{y_{1}y_{0},m_{1}}(t,m_{0})\leq 1$
$\sum_{y_{1}=0}^{1}\sum_{y_{0}=0}^{1}\sum_{m_{1}=0}^{1}\psi_{y_{1}y_{0},m_{1}}(t,m_{0})=1.$
Let $W_{i}=(Y_{i},M_{i},T_{i},Z_{i})$ be the observed data vector with possible realized value $w\equiv(w_{y},w_{m},w_{t},w_{z})$. Abusing notation, we let $w_{1}=(w_{y},w_{m},w_{t},1)$ and $w_{0}=(w_{y},w_{m},w_{t},0)$. The randomization and consistency assumptions imply the following moment conditions:
$\mathbb{E}[g_{w_{1}}(W_{i},\psi)\mid Z_{i}=1]$ $=\mathbb{P}(Y_{i}=w_{y},M_{i}=w_{m},T_{i}=w_{t}\mid Z_{i}=1)$
$\qquad-Q_{*}\left(\sum_{y_{0}}\psi_{w_{y}y_{0}w_{m}}(w_{t},1)\right)-(1-Q_{*})\left(\sum_{y_{0}}\psi_{w_{y}y_{0}w_{m}}(w_{t},0)\right)$
$\mathbb{E}[g_{w_{0}}(W_{i},\psi)\mid Z_{i}=0]$ $=\mathbb{P}(Y_{i}=w_{y},T_{i}=w_{t}\mid M_{i}=w_{m},Z_{i}=0)-\left(\sum_{y_{1}}\sum_{m_{1}}\psi_{y_{1}w_{y}m_{1}}(w_{t},w_{m})\right).$
There are $d_{1}=8$ for the post-test data and $d_{0}=8$ restrictions for the pre-test data.
Define $r_{d}(\psi)$ encode the deterministic restrictions on the $\psi$ values and let $\Psi^{\dagger}_{d}=\{\psi\;:\;r_{d}(\psi)\geq 0\}$ be the values of the underlying parameters that satisfy these restrictions. These restrictions include assumptions like monotonicity that would cause certain $\psi$ values to be set to zero or the sensitivity analysis specifications that limit the size of a group of $\psi$ values. Under these maintained assumptions, we can characterize the identified set as
$\Psi^{\star}=\{\psi\in\Psi^{\dagger}_{d}\;:\;\mathbb{E}[g_{w_{1}}(W_{i},\psi)]=0,E[g_{w_{0}}(W_{i},\psi)]=0\;\forall w_{1},w_{0}\in\{0,1\}^{3}\}.$
One way to define a distance from the identified set is with a population criterion function. Let the moment conditions be indexed by $j$ such that $j=1,\ldots,8$ correspond to $\{g_{w_{1}}\}$ and $j=9,\ldots,16$ correspond to $\{g_{w_{0}}\}$. Then, the population criterion (loss) function is
$L(\psi)=\sum_{j=1}^{16}|\mathbb{E}[g_{j}(W_{i},\psi)]|$ (14)
Following Torgovitsky (2019), we use absolute value loss here to ensure that we can leverage linear programming techniques for computational convenience. We can obtain an empirical version of the criterion,
$L_{n}(\psi)=\sum_{j=1}^{16}\sqrt{n}|\overline{g}_{j}(W_{i},\psi)|,$ (15)
where, for instance,
$\overline{g}_{w_{1}}(W_{i},\psi)$ $=\mathbb{P}_{n}(Y_{i}=w_{y},M_{i}=w_{m},T_{i}=w_{t}\mid Z_{i}=1)$
$\qquad-\widehat{Q}_{*}\left(\sum_{y_{0}}\psi_{w_{y}y_{0}w_{m}}(w_{t},1)\right)-(1-\widehat{Q}_{*})\left(\sum_{y_{0}}\psi_{w_{y}y_{0}w_{m}}(w_{t},0)\right),$
and $\mathbb{P}_{n}$ is the in-sample distribution.
We could proceed with estimating the bounds for a given set of assumptions by searching over the parameter space such that $L_{n}=0$, but this is often a fragile approach. In particular, it may be the case that restrictions hold in the population but fail to hold in empirical samples due to sampling variability so that the minimum value of $L_{n}$ is strictly greater than 0. As an alternative approach, we can first find the minimum value of $L_{n}$ in the sample and then find extreme values of the parameter under parameter values that are close to that minimum.
We first define the sample minimum of the criterion function under the maintained deterministic restrictions,
$\overline{L}_{n}=\inf_{\psi\in\Psi^{\dagger}_{d}}L_{n}(\psi).$ (16)
We can then estimate the upper and lower bounds by finding the minimum and maximum values of $\delta$ that come close to this value:
$\widehat{\delta}_{L}$ $=\min_{\psi\in\Psi^{\dagger}_{d}}\delta(\psi)\;\text{s.t.}\;L_{n}(\psi)\leq\overline{L}_{n}(1+\epsilon_{n}),$
$\widehat{\delta}_{U}$ $=\max_{\psi\in\Psi^{\dagger}_{d}}\delta(\psi)\;\text{s.t.}\;L_{n}(\psi)\leq\overline{L}_{n}(1+\epsilon_{n}).$
The tuning parameter $\epsilon_{n}$ controls how close we require the criterion function of the bounds to be to the overall sample minimizer. This approach requires $\epsilon_{n}\to 0$ as $n\to\infty$. We take $\epsilon_{n}=0.25$ in our implementation, which has shown good performance in simulations.
For the confidence intervals, we use the nonparametric bootstrap to obtain standard error estimates of each of the bounds and then apply the (Imbens and Manski, 2004) approach. In simulations, we found this approach to be slightly conservative and other competing methods, such as Chernozhukov, Newey and Santos (2023), to undercover the true parameter slightly, at least in simulations similar to our data example. Thus, we use the nonparametric bootstrap approach for our confidence intervals.
Appendix C Parametric Bayesian Approach to Incorporate Covariates
The nonparametric bounds above are sharp in the sense that they leverage all information about the outcome, moderator, treatment, and question order. Researchers, however, often have additional data in the form of covariates that may help reduce the uncertainty of their estimates. Here, we consider a Bayesian parametric model of the principal strata approach to the pre-test, post-test, and random placement designs, building on the work of Mealli and Pacini (2013) (see also Imbens and Rubin, 1997; Hirano et al., 2000). Unlike the nonparametric bounds approach, a Bayesian model allows us to incorporate prior information about the data-generating process in a smooth and flexible manner. [^4] Levis et al. (2023) proposes a way to incorporate covariates on nonparametric bounds when the quantity of interest can be written as an average of covariate-specific quantities. Unfortunately, we cannot write the interaction in this way because it is the difference between two different CATEs that condition on different subsets of the data. One could use their approach on each of the individual CATEs and combine those bounds for the interaction, but the resulting bounds would not be sharp.
C.1 The Model
Our approach focuses on a data augmentation strategy that models the joint distribution of the outcomes and the principal strata, the latter of which are not directly observable. We allow the distribution of the potential outcomes and principal strata conditional on those strata to further depend on covariates via a binomial and multinomial logistic model, respectively:
[…]
Our goal is to make inferences about the posterior distribution of these parameters and the ultimate quantities of interest, $\tau(d)$ and $\delta$. There are two ways to represent these quantities under this parametric model, resulting in two different types of posterior distributions. The first is based on population inference and derives expressions for $\tau(d)$ and $\delta$ purely in terms of the parameters of the model. The second is based on in-sample inference and derives expressions for $\tau(d)$ and $\delta$ in terms of potential outcomes in a particular sample.
[…]
$\tau_{i}(m)=\mathbb{E}\left[Y^{*}_{i}(1)-Y^{*}_{i}(0)\mid M^{*}_{i}=m,X_{i}\right]\ =\ \sum_{s\in\mathcal{S}^{*}_{d}}(\mu_{is}(1,1)-\mu_{is}(0,1))\rho_{is},$
[…]
The in-sample versions of the quantities of interest are more straightforward, since they are simply the conditional mean differences and interaction among the units in the sample,
$\tau_{s}(m)=\frac{\sum_{i=1}^{n}\mathbb{I}(M^{*}_{i}=m)\left\{Y^{*}_{i}(1)-Y^{*}_{i}(0)\right\}}{\sum_{i=1}^{n}\mathbb{I}(M^{*}_{i}=m)},$
and $\delta_{s}=\tau_{s}(1)-\tau_{s}(0)$. Obviously, across repeated samples, we can relate these to the population quantities as $\mathbb{E}[\tau_{s}(m)]=\tau(m)$ and $\mathbb{E}[\delta_{s}]=\delta$.
We develop an efficient Markov Chain Monte Carlo (MCMC) algorithm to take draws from the posterior and then calculate these quantities of interest. Our Gibbs sampler can also be simplified and used for inference in the absence of pre-treatment covariates, which can be viewed as a Bayesian alternative to uncertainty estimation for the partially identified parameters discussed in Section 4.4. We provided details of these algorithms in Supplemental Materials D.
This Bayesian approach has the advantage of easily incorporating covariates, but it does require us to select prior distribution for the model parameters, some of which are unidentified in the frequentist sense. Thus, the identification of these parameters will depend on the prior. To investigate this, we take draws of the prior predictive distribution under different prior structures, which we show in Figure SM.6. All the priors we considered are symmetric, but uniform priors on the model parameters lead to somewhat informative priors on the ultimate quantity of interest. Thus, we rely on more dispersed priors for the simulations and the application. We discuss the choice of prior distribution more fully in Supplemental Materials D.
We conduct two simulation studies to demonstrate the gains in efficiency from the monotonicity and stability assumptions and the incorporation of covariates in the Bayesian approach. The first simulation varies the assumptions of the data-generating process and compares the posterior variance of these distributions across combinations of our two assumptions. The second simulation varies the predictive power of the covariates on the outcome and the strata in the data-generating process and compares the variance of the posterior distributions from Gibbs run on each simulated data set with and without incorporating covariates. We present these results in Supplemental Materials  D.1.
Appendix D MCMC Algorithm
In this section we describe our MCMC algorithm for the Bayesian model of Section C. Our goal is to sample from the joint distribution of the parameters and the principal strata indicator,
[…]
Conditional on the other parameters, then the full conditional posterior of the principal strata follows a similar form to Hirano et al. (2000),
[…]
where we suppress the dependence of $\mu_{is}$ and $\rho_{is}$ on the model parameters. Repeatedly drawing from these full conditional posterior distributions should provide a sample from the above joint posterior and allow for posterior inference in the usual manner. In each iteration, $r\in\{1,\ldots,R\}$, of the algorithm, we have draws
[…]
We can use these draws to generate draws of the population and in-sample versions of the quantity of interest. Given that $\widehat{S}^{(r)}_{i}$ is the imputed principal strata imputed for unit $i$ in the $r$th draw from the posterior, we let
[…]
be the mean of the potential outcomes conditional on that imputed principal strata. Furthermore, let $\widehat{\rho}^{(r)}_{is}$ be the $r$th draw of the predicted probabilities of each principal strata for each unit. Then, we can calculate the population quantity as
$\widehat{\delta}_{p}^{(r)}=\frac{1}{n}\sum_{i=1}^{n}\left(\sum_{s\in\mathcal{S}^{*}_{1}}(\widehat{\mu}^{(r)}_{is}(1,1)-\widehat{\mu}^{(r)}_{is}(0,1))\widehat{\rho}^{(r)}_{is}\right)-\left(\sum_{s\in\mathcal{S}^{*}_{0}}(\widehat{\mu}^{(r)}_{is}(1,1)-\widehat{\mu}^{(r)}_{is}(0,1))\widehat{\rho}^{(r)}_{is}\right),$
For the in-sample quantity, we can then draw imputed values of the missing potential outcomes themselves $\widehat{Y}^{*,(r)}_{i}(1)\sim\text{Bin}(\widehat{\mu}^{(r)}_{i}(1,1))$ and $\widehat{Y}^{*,(r)}_{i}(0)\sim\text{Bin}(\widehat{\mu}^{(r)}_{i}(0,1))$. We can combine this with the imputed value of $M^{*}_{i}$, which mechanically derives from $\widehat{S}_{i}^{{r}}$, to get the $r$th draw from the posterior of $\delta_{s}$,
$\widehat{\delta}^{(r)}_{s}=\frac{\sum_{i=1}^{n}\widehat{M}^{*,(r)}_{i}\left\{\widehat{Y}^{*,(r)}_{i}(1)-\widehat{Y}^{*,(r)}_{i}(0)\right\}}{\sum_{i=1}^{n}\widehat{M}^{*,(r)}_{i}}-\frac{\sum_{i=1}^{n}\left(1-\widehat{M}^{*,(r)}_{i}\right)\left\{\widehat{Y}^{*,(r)}_{i}(1)-\widehat{Y}^{*,(r)}_{i}(0)\right\}}{\sum_{i=1}^{n}\left(1-\widehat{M}^{*,(r)}_{i}\right)}.$
Broadly speaking, we would not expect very large differences between these two targets, except for slightly less posterior variance for the in-sample version.
Figure SM.6: Prior predictive distribution of the parameter under three different prior distributions: (red) the default priors that scales a Jeffreys prior by the number of principal strata; (blue) a uniform prior on all parameters; and (green) a more extreme prior that has $\alpha=1/16$.
As discussed in the main text, the priors need careful attention because they drive the identification of the parameters that are unidentified by the likelihood. One additional complication comes from how the ultimate quantity of interest is a function of the parameters so we cannot directly place, for example, a uniform prior on $\delta$. Figure SM.6 shows the prior predictive distribution for interaction with three different priors when monotonicity and stable moderator under control are assumed and there are no covariates. The uniform prior on all parameters results in a prior on $\delta$ that has more density in the center of identified range than we might expect. This result is similar to how sums of uniform random variables are not themselves uniform. We can counteract this issue by reducing the Dirichlet and Beta hyperparameters below 1 to put more density at extreme values of the parameters compared to the center. Dropping these parameters down to 1/16 (in green) leads to more mass on strata means closer to 0 or 1 and strata probabilities closer to 0 and 1. In terms of the interaction, this leads to more mass at the values -2, -1, 0, 1, and 2. Our default prior (red) is one that scales the hyperparameters by the inverse of the number of strata to achieve something closer to a uniform distribution.
Additionally, we re-ran the Gibbs empirical analysis of the Horowitz and Klaus (2020) study, adjusting the priors to the extreme values or uniform values in the previous simulation. The results are displayed in Figure SM.7, demonstrating the general consistency of the point estimates across starting priors, although there is some fluctuations in the variance of the results.
Figure SM.7: Comparing Bayesian estimates for $\delta$ for default, extreme, and uniform priors
Notes: Figure shows posterior means and 95% credible intervals for $\delta$ under different sets of assumptions, applied to either the post-only data (grey) or the combined pre-post data (black). Estimates are shown with default, extreme, and uniform priors, (denoted by circles, triangles, and squares, respectively). We follow the original authors in using age, gender, education, and closeness to one’s ethnic group as covariates. The naïve OLS estimates are included for comparison.
D.1 Simulation Evidence for the Bayesian Approach
Simulation Study I.
In the first simulation study, we generate simulated data with $n=1000$ constructed using a data generating process that matches the Bayesian posterior, pre-specifying coefficient values for the outcome and principal strata models, randomly drawing values of $Z$, $T$, and three covariates $X_{1}$, $X_{2}$, and $X_{3}$, and generating values of $Y$ and $M$ from the models. Tables SM.2 and SM.3 in the Supplemental Materials display the $\beta$ coefficients for the outcome and $\psi$ coefficients for the for the true data generating process (DGP). The DGP assumes that monotonicity and stable moderator under control both hold so that there are three feasible strata ($\mathcal{S}=\{000,100,111\}$). Thus, in this setting it would be most appropriate to incorporate both assumptions into the MCMC algorithm for sampling from the posterior distribution. Since these assumptions narrow the nonparametric bounds, we expect the assumptions to reduce variance of the posterior distribution of $\delta$.
To test this, we perform a Monte Carlo simulation with 1,000 iterations. For each iteration, we calculate the posterior distribution of $\delta$ with the same data across four different versions of the MCMC algorithm: enforcing just the monotonicity assumption, enforcing just the stable moderator under control assumption, enforcing neither assumption, and enforcing both assumptions. Each run of our MCMC algorithm consists of 4 chains with 2,000 iterations each, 200 burn-in (or warm-up) iterations, and a thinning parameter of 2. Both in-sample and population $\delta$ values are calculated at each iteration and the variance of the posterior is calculated from a sample of 1,000 draws from the posterior. This is done for each of the 1,000 simulated datasets, and for each dataset we compute the percent reduction in variance compared to the MCMC algorithm with no assumptions when using the algorithm with the monotonicity assumption, the stable moderator under control assumption, or both assumptions.
Figure SM.8: Variance reduction from different combinations of assumptions. Boxplots present distribution of % variance reduction of $\delta$ from the MCMC algorithm with the labeled assumptions compared to same algorithm with no assumptions, across 1,000 draws of simulated data. For each simulation iteration, 4 chains were run for each combination of assumptions. MCMC parameters: 2,000 iterations, 200 burn-in, 2 thinning parameter, simulated data $n=1,000$.
Figure SM.8 presents boxplots for the distribution of reductions in variance for each combination of assumptions. Both the monotonicity and stable moderator assumptions on their own reduce the variance compared to no assumptions, while making both assumptions reduces the variance even further. The monotonicity assumption showed a median posterior variance reduction of 40.8% for the in-sample $\delta$ and 42.0% for the population $\delta$. The stable moderator under control assumption on reduced the posterior variance by a median reduction of 47.1% (in-sample $\delta$) and 50.4% (population $\delta$). The MCMC algorithm with both assumptions exhibited a posterior variance reduction of 59.3% (in-sample $\delta$) and 61.7% (population $\delta$).
Simulation Study II.
In the second simulation study, we drew a series of simulated datasets under different conditions where the covariates had a weak, medium, or strong correspondence with the outcome and principal strata in the data generating processes. Thus, there were six total conditions: Weak, Medium, and Strong influence in the outcome DGP; and Weak, Medium, and Strong influence in the principal strata DGP. The values of the coefficients for these conditions are $\beta$ and $\psi$ values of 0, 0.25, and 0.5, respectively. When varying the influence of covariates in the outcome DGP, the influence of covariates on the strata was held constant, and the influence in the outcome model was similarly held constant when varying influence in the strata DGP. Fixed values of the $\beta$’s and $\psi$’s are shown in the Supplemental Materials in Tables SM.4 and SM.5.
Figure SM.9: Variance reduction from incorporation of covariates. Boxplots present distribution of % variance reduction of $\delta$ from Gibbs with covariates compared to MCMC without covariates, across 1,000 draws of simulated data. For each draw 4 MCMC chains were run for each combination of assumptions. MCMC parameters: 2,000 iterations, 200 burn-in, 2 thinning parameter, simulated data $n=1,000$.
For each condition, we drew 1,000 simulated datasets and ran the MCMC algorithm twice: one time incorporating covariates and one time omitting them. Each MCMC run consisted of the same iterations, burn-in, and thinning parameters as in the previous simulation study. We again calculate in-sample and population $\delta$ values for each iteration of the Gibbs and calculate the variance of the posterior distribution and the % variance reduction comparing the Gibbs with covariates to that without. Figure SM.9 presents boxplots for the distribution in variance reduction. When we vary the influence of covariates on the outcome, we see a clear variance reduction in all conditions, and we observe a larger reduction as the influence of covariates on the outcome in the DGP increases. When testing the impact of incorporating covariates across different levels of influence in the DGP on the strata, the pattern is less pronounced, with overall reduction increases in all conditions but slightly lower reductions in the Medium than Weak condition. The Strong condition still has the largest variance reduction overall, however, so in general the efficiency gains from incorporating covariates are increasing as the influence of covariates on strata in the data increases.
Appendix E Additional Simulation Details
Table SM.2: $\beta$ Values for DGP in Bayesian Assumptions Simulation
Variable $\beta$
(Intercept) -2.00
X1 1.00
X2 0.15
X3 (Medium) 0.24
X3 (Large) 0.28
T 0.83
Z -0.01
T:Z 0.11
S111 0.41
S100 0.62
T:S111 0.01
T:S100 0.23
Z:S111 0.20
Z:S100 -0.02
T:Z:S111 -0.90
T:Z:S100 0.09
Table SM.3: $\psi$ Values for DGP in Bayesian Assumptions Simulation
S111 s100 s000
(Intercept) -2.06 -1.00 0.00
X1 2.00 1.50 0.00
X2 0.50 0.17 0.00
X3 (Medium) 1.35 -0.28 0.00
X3 (Large) 1.75 -1.01 0.00
Table SM.4: Fixed $\beta$ Values in Covariate Simulation
Variable $\beta$
(Intercept) -1.00
X1 1.00
X2 0.50
X3 (Medium) 0.50
X3 (Large) 0.28
T 0.83
Z -0.01
T:Z 0.11
S111 0.41
S100 0.62
T:S111 2.00
T:S100 -0.13
Z:S111 0.50
Z:S100 0.10
T:Z:S111 0.05
T:Z:S100 0.01
Table SM.5: Fixed $\psi$ values in Bayesian Covariate Simulation
S111 S100 S000
(Intercept) -2.06 -1.00 0.00
X1 2.00 1.50 0.00
X2 0.50 0.17 0.00
X3 (Medium) 1.35 -0.28 0.00
X3 (Large) 1.75 -1.01 0.00
E.1 Comparing the sharp bounds and Bayesian approach without covariates
Figure SM.10: Comparing non-parametric bounds and Bayesian estimates for $\delta$ under different assumptions
Notes: The figure shows nonparametric bounds and Bayesian estimates of the quantity of interest under different sets of assumptions applied to either the post-only data (grey) or the combined pre-post data (black). The thick bars denote the width of the bounds, and thinner lines denote the 95% confidence intervals around the bounds. Across the first four panels, the thin lines with dots denote the Bayesian posterior mean and 95% credible interval. This estimate included no covariates to facilitate comparison with the nonparametric bounds. For the final panel (“OLS”), the thin lines with dots denote the OLS estimate and 95% confidence interval.
Figure SM.10 displays the non-parametric bounds (with 95% confidence intervals) and Bayesian estimates (posterior means with 95% credible intervals) for $\delta$ under different sets of assumptions applied to the post-only data (in grey) and to the combined pre-post data (in black). We also include the naïve OLS estimate with 95% confidence interval for comparison in the final panel.
E.2 Incorporating covariates into the Bayesian approach
Figure SM.11: Comparing Bayesian estimates for $\delta$ with and without covariates
Notes: Figure shows posterior means and 95% credible intervals for $\delta$ under different sets of assumptions applied to either the post-only data (grey) or the combined pre-post data (black). Estimates are shown with and without the inclusion of covariates (denoted by triangles and circles, respectively), and the numbers indicate the reduction in the width of the credible intervals due to the inclusion of covariates for the post-only data (in grey) and the combined pre-post data (in black). We follow the original authors in using age, gender, education, and closeness to one’s ethnic group as covariates. The naïve OLS estimates are included for comparison.
Thus far, our Bayesian estimates have omitted covariates to aid comparison with the nonparametric bounds. However, as discussed above, a key attraction of the Bayesian approach is the ease with which we can incorporate additional information. Figure SM.11 presents posterior means with 95% credible intervals, both with and without covariates, under different assumptions. We follow the original authors in using age, gender, education, and closeness to one’s ethnic group as covariates. Including covariates significantly tightens the credible intervals, especially when fewer assumptions are imposed. For example, when only randomization is assumed, the width of the 95% credible interval shrinks by more than 50% when including covariates. While including covariates does not alter our substantive conclusions in this case, it does show that incorporating additional information can lead to large gains in precision. Since researchers often include a wide range of control variables in the design of a survey experiment, flexibly leveraging this information is a key advantage of the Bayesian approach.